OpenStudy (anonymous):

Factor the polynomial: 2a^4 - 9a^2 + 4

jagr2713 (jagr2713):

is there an answer choice

OpenStudy (anonymous):

No

OpenStudy (anonymous):

since the degree of 1st term is just twice the degree of the 2nd term and 3rd term is a constant, we can make use of some substitution such as a^2=x then re-write the equation as\[2x^2-9x+4=0\]which is easier to solve, simply a quadratic equation... solve for x this time...

OpenStudy (anonymous):

after solving for x... you can solve for a

OpenStudy (anonymous):

can you do it from here? @yankeeez

OpenStudy (anonymous):

ok... factoring...\[2x^2-9x+4=0\]\[(2x-1)(x-4)=0\]therefore\[x=1/2~and~4\]solving for \(a^2=x\), if x=1/2...\[a=\pm\sqrt{1/2}=\pm\frac{\sqrt{2}}{2}\], and if x=4...\[a=\pm\sqrt4=\pm2\]therefore the roots of \(2a^4-9a^2+4=0\) are \(\pm2,\pm\sqrt2/2\).